Income elasticity of demand Calculate the income elasticity of demand for the following demand functions and...

  1. Income elasticity of demand

Calculate the income elasticity of demand for the following demand functions and assess its magnitude (is the good normal, inferior, a luxury or a necessity?).

  1. Q =50 -10p + 3Y, at p = 2 and Y = 20

  1. Q = 5 - 20p + 4Y, at p = 1 and Y = 50
  1. Q = 70 – 3p – 2Y, at p = 15 and Y = 5

  1.   Q = 35 - 20p + 0.5Y, at p = 1.5 and Y = $1,000,000,000
Answer #1

Income elasticity of demand, E =

where dQ/dY can be found using demand function.

a) Q = 50 -10p + 3Y, at p = 2 and Y = 20
So, Q = 50 - 10(2) + 3(20) = 50 - 20 + 60 = 90
E = (3)*(20/90) = 0.67
As 0 < E < 1 so the good is normal good and it is a necessity.

b) Q = 5 - 20p + 4Y, at p = 1 and Y = 50
So, Q = 5 - 20(1) + 4(50) = 5 - 20 + 200 = 185
So, E = 4*(50/185) = 1.08
As E > 1 so the good is normal good and it is a luxury.

c) Q = 70 – 3p – 2Y, at p = 15 and Y = 5
So, Q = 70 – 3(15) – 2(5) = 70 - 45 - 10 = 15
So, E = (-2)*(5/15) = -0.67
As E < 0, so the good is inferior.

d) Q = 35 - 20p + 0.5Y, at p = 1.5 and Y = $1,000,000,000
So, Q = 35 - 20(1.5) + 0.5(1,000,000,000) = 35 - 30 + 500,000,000 = 500,000,005
So, E = (0.5)*(1,000,000,000/500,000,005) = 1
As E = 1, so the good is normal.